OpenStudy (anonymous):

help please.. Stock A B C Price f(x) = 25(1.08)x x f(x) 0 22.00 1 21.56 2 21.13 The initial price of this stock is $30, but it has been increasing 4% each day. Create the price function for Stock D. It should have the second lowest starting price and the second highest rate of growth. Using complete sentences, justify that your function meets these requirements.

OpenStudy (anonymous):

please help me too once ur done here:) i would appreciate it guys

OpenStudy (anonymous):

Graph the price function for Stock D. Explain the key features of the graph with complete sentences. A recent drop in sales has affected Stock D with the function g(x) = –6. Explain to Gordon how Stock D’s new price function, f(x) + g(x), will be created. Graph f(x) + g(x). Gordon wants to fully understand what kind of changes could affect the money he is investing. Using complete sentences, explain the effect of the following transformations. Graph them and f(x) on the same coordinate plane. f(x + 2) f(–x) f(x) + 3

OpenStudy (tkhunny):

That's not very clear. Let's see if I can sort it out. Stock A: \(f(x) = 25\cdot 1.08^x\) Stock B: \(f(0) = 22.00,\;f(1)=21.56,\;and\;f(2) = 21.13\) Stock C: \(f(x) = 30\cdot 1.04^x\) Starting Price: A: 25, B: 22.00, C: 30 Growth Rate: A: 8%, B: -2.00%, C: 4% Your task is to model Stock D with: Starting Price: (22.00, 25) Growth Rate: (-2%, 4%) Well, do it! There are infinitely many to choose from.

OpenStudy (anonymous):

idk how to figure this out can u solve it? @tkhunny

OpenStudy (tkhunny):

Yes, I can. But that won't help you learn. Stock B can be rewritten as \(f(x) = 22\cdot 0.98^x\) There you have three fine examples. Just invent one that meets your criteria. You have to perform the last requirement. Write and justify!

OpenStudy (anonymous):

can you just awnser it please i need this assigment done so i can finish my module and take a test @tkhunny

OpenStudy (anonymous):

i could help u @lovelychick13

OpenStudy (anonymous):

i could help u @lovelychick13. wht u need help wit?

OpenStudy (anonymous):

im good now thanx

OpenStudy (anonymous):

oh:(

OpenStudy (tkhunny):

Absolutely I could just do it. But, I cannot not do it FOR YOU. I can do it only in place of you. The task, here, is for you to learn. If you are not willing to try, I cannot help you. If it makes no sense at all to you, you may be in the wrong class. That's a different problem.